geerky42 (geerky42):

Find the area of largest isosceles triangle that can be inscribed in a circle of radius 4.

geerky42 (geerky42):

I have to use calculus to solve this problem. Where should I start?

OpenStudy (abb0t):

I had a problem similar 2 this on my old exam! ahh! Give me a minute haha. I know this!! Lol

OpenStudy (abb0t):

Let Q be the apex angle. Draw the altitude of the triangle (bisecting the apex angle) through the center of the circle. This altitude has length 4 + h, where h = 4 cos(Q) and r = 4 Base of triangle has length: 8 sin(Q) So the area is: \[A = \frac{ 1 }{ 2 }(4+4\cos(Q)(8\sin(Q) = 24\sin(Q) + 12\sin(2Q)\] \[\frac{ dA }{ dQ } = 24cos(Q) + 24 cos(2Q)\] Set this expression to zero to find the optimal triangle. \[\cos(Q)+2\cos^2(Q)-1=0\] (2 cos(Q) - 1)(cos (Q) + 1 ) = 0

geerky42 (geerky42):

I don't get it, why is altitude 4 + h?

OpenStudy (abb0t):

the height=r+h

geerky42 (geerky42):

I'm sorry, but can you sketch it for me?

geerky42 (geerky42):

"height=r+h" What does h represent?

OpenStudy (abb0t):



OpenStudy (abb0t):

I used So, b=cos(theta) and h=rsin(theta)

geerky42 (geerky42):

Ah, I see. Thanks.

OpenStudy (abb0t):

Sorry, this type of problem is hard to explain. I am using an my old exam to figure out how to solve this. I will definitely post more if I can find a better way to explain.

OpenStudy (abb0t):

But I used area as a function of half the angle

geerky42 (geerky42):

Ok, thank you, I appreciate it.

OpenStudy (abb0t):

so the area is \[A = \frac{ 1 }{ 2 }b(r+h)\] Substituted for b and h (r=4) and get A as a function of theta, so then I could solve the problem by differentiating. Also, I used b=cos(theta) and h=rsin(theta)

OpenStudy (anonymous):

@geerky42 My answer is A = 12√3 unit² My way is to take advantage of the symmetry characteristic between circle and isosceles triangle. Let me know if you want further details :)